OpenStudy (anonymous):

which property is illustrated by the following? 8x-17=51 8x=68

myininaya (myininaya):

addition b/c 51+17=68

OpenStudy (anonymous):

associative because you can do this:

myininaya (myininaya):

associative is (a+b)+c=a+(b+c)

myininaya (myininaya):

or (ab)c=a(bc)

myininaya (myininaya):

8x-17+17=51+17 8x=68

OpenStudy (anonymous):

Sorry I was trying to split it up and show it could be the same on both sides

myininaya (myininaya):

or maybe its inverse property on one side and addition property on the other side

OpenStudy (anonymous):

8x-17-51=0 same as 8x-68=0

OpenStudy (anonymous):

a-b-c (a-b)-c NOT the same thing as a-(b-c) so not associative.

OpenStudy (anonymous):

Additive is not a property though: http://mathforum.org/dr.math/faq/faq.property.glossary.html

OpenStudy (anonymous):

You know what this is? It's the property of equality. They are equal.

OpenStudy (anonymous):

8x-17=51 is simply equivalent to: 8x=68 Therefore it's an illustration of the property of equality.

OpenStudy (anonymous):

I at first thought it could be associative because you can break 68 up into 51+17, but it is not correct.

OpenStudy (anonymous):

To be clear, this is: The symmetric property of equality, which says that if a = b, then b = a.